OpenStudy (anonymous):

what number must be added to complete the square x^2+7x

hartnn (hartnn):

hi :) \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) since we know \(\large (a+b)^2 = a^2+2ab+b^2\) we have 2b as 7 here and we need b^2

hartnn (hartnn):

so from 2b = 7 first find b and then square it, thats it!

OpenStudy (anonymous):

b=7/2 b=(7/2)^2